# Upgrading Democracies with Fairer Voting Methods

### v1

@@ 3 Discussion
Fair voting methods to foster a novel impact creation process. The influence of the chosen voting method on the proposed and selected project ideas can transform the process of creating real-world impact. For instance, novel welfare projects gain momentum, as they appeal to several interest groups, while they may also be easier to implement than costly infrastructural projects [27]. We foresee that, on the long term, proportional voting methods could enable an incremental grass-roots implementation of sustainable development projects over several participatory budgeting campaigns, which could also support a fairer transition to a net zero future.
@@
Fair voting methods: a better-than-optimal paradigm. Over-simplifying, averaging or overriding people’s preferences while optimizing societies for goals that people may not even explicitly agree on is a significant risk that democracies face in the era of artificial intelligence (AI) run by Big Tech. In this context, the notion of ‘optimal’ solutions is actually misleading and displaced from the reality of people, as demonstrated by the inferiority of popular expensive projects or projects focused on city centers that are typically proposed in participatory budgeting processes under majoritarian standard voting. In contrast, proportional voting methods work well for pluralistic settings, where voters may not agree on one and the same goal, but decisions are based on collective intelligence. This approach excels when goals are complementary, but is also suited to handle situations of competitive or even incompatible goals, or when significant minorities are involved. Small compromises, such as excluding a popular expensive project, open up opportunities for large collective returns with benefits for a great majority of people. For instance, the inclusion of under-represented health projects or of projects in the city periphery can reduce social tensions and benefit a broader population. The demonstrated approach of combining cumulative voting with the method of equal shares delivers a powerful blueprint for upgrading democracies. It allows societies to achieve (i) a larger number of realized projects within the respectively available budget; (ii) a greater resourcefulness of novel projects and a multi-faceted impact; (iii) more participation and inclusion; (iv) greater transparency and trust in the voting process; (v) better representation of citizen preferences and higher satisfaction; (vi) more (spatial) fairness; (vii) higher legitimacy thanks to a greater acceptance of the collective outcomes; (viii) strengthened democratic values of voters. We refer these impacts "better than optimal" as they tend to be a win-win for all, while they are distinguished from what existing optimization and data-driven decision-making approaches typically offer.
@@
Fair voting methods in the era of artificial intelligence. We would like to end on a remark addressing the use of data-driven AI systems for decision making. While we consider AI a very powerful technology with numerous promising application areas, how to use it in the context of governing democracies must be studied very carefully. AI safeguards for democracy are critical and yet to be created. Existing decision-making methods supported by AI solutions show alarming vulnerabilities and are still perceived as lacking fairness and legitimacy [39, 40, 41]. For instance, biases and inconsistencies by unconsolidated use of large language models can exacerbate by voting methods with a deficit on fairness [42]. It is not clear when and to what extent AI risks can be addressed. Therefore, alternative approaches such as the one presented here are urgently needed.
@@ 4 Methods
We illustrate the designed participatory budgeting campaign of City Idea, the evaluation approach of legitimacy and the models for causal explanations for the shifts between ballot aggregation methods before and after voting.
4.1 The participatory budgeting campaign of City Idea: field test design
Table 1 outlines the participatory budgeting campaign of City Idea. It consists of six phases over the passage of nine months. An entry and exit survey were conducted before and after voting respectively. The entry survey was launched via physical invitation letters to all citizens sent by city council (10,000 A5 flyers [29]). This resulted in 3592 respondents. Some of these filled in their email to voluntarily participate in the exit survey and this resulted in 808 respondents. The surveys included questions about demographics and political opinion (Table LABEL:table:entry-demographics-LABEL:table:exit-demographics), engagement and digital skills (Table LABEL:table:entry-digital-LABEL:table:exit-digital), the City Idea voting (Table LABEL:table:entry-city_idea-LABEL:table:exit-city_idea), the voting outcome preference (Table LABEL:table:entry-voting_outcome-LABEL:table:exit-voting_outcome) and the voting input and aggregation method (Table LABEL:table:entry-voting_input-LABEL:table:exit-voting_input). Survey questions encode variables, which are observed for changes before and after voting. The overall study has received ethical approval from University of Fribourg (#2021-680).
Table 1: The rigorously designed participatory budgeting campaign of City Idea with 6 phases over a period of 9 months in 2023. Participants of the entry survey, voting and exit survey are linked using an anonymous identifier (hash) of the citizens’ social security number. This allows the study of several choice variables before and after voting.
@@ 6. Exit survey July September 808 participants
Project ideation, deliberation and campaigning. Citizens proposed more than 161 project ideas out of which 33 projects are selected to put for voting [29]. Projects are eliminated or similar project ideas are combined to strengthen their chances to win. Project ideation involved support and guidance with planning, implementation, economic and legal feasibility by a municipal advisory board of 7 people, a core team of 3 people and ambassadors. The city council distributed posters and flyers to various municipal institutions and locations. It also organized two press releases that resulted in 17 media reports. Moreover, the interactive information campaign ‘know your projects’ was launched in the city that engaged 73 participants in 1682 random pairwise comparisons of projects (see Table LABEL:tab:pairwise_winners for further information) via the Stanford Participatory Budgeting platform  [43]. Awareness about the method of equal shares and what projects are expected to win was raised, emphasizing the challenge of electing projects that cover a large share of the budget. Section LABEL:sec:equal-shares-explanation provides more information about how the method of equal shares was communicated and explained to citizens.
Voting. The city of Aarau has a total of 22,593 citizens, out of which 17,428 are Swiss citizens. A total of 1703 voters participated in the City Idea campaign, out of which 1330 are citizens of Aarau. This also allows individuals to have a voice, who are not living in Aarau, but have a connection to the city, such as workers and broader family members. Children over 15 years old could also vote. Voting was digital, using the open-source Stanford Participatory Budgeting platform [43]. However, voters with low digital literacy were supported in virtual voting centers around the city. The voting platform was extended to support cumulative voting by distributing 10 points over at least 3 projects [30, 43]. Voters are securely verified in a privacy-preserving way by providing a unique (hash) code that was generated by the city council using the social security number of each citizen (AHV card). The same platform was used for project campaigning by using the pairwise project comparisons feature. The winning projects were calculated using the method of equal shares and the Add1U completion method implemented within the Pabutools (the ones of Świecie and Wieliczka used the Add1 completion method, while Add1U would yield 1 and 3 more winners) [37]. Voting outcomes with the utilitarian completion method were also calculated and assessed in the exit survey (Table LABEL:table:exit-voting_outcome).
@@ 4.2 Evaluation of legitimacy
To assess the legitimacy of ballot aggregation methods, we rely on well-established theory of political legitimacy [34] and social choice [44, 45]. Among the three notions of input (voting participation), throughput (the voting process) and output (voting outcome) legitimacy [46, 47], we mainly focus in Section 2.2 on output legitimacy. However, we also assess the choice of ballot aggregation method before and after voting with equal shares as well as the effect of explaining equal shares, both capturing throughput legitimacy.
As such, by legitimacy we refer to voters who remain satisfied and accept the voting outcome of a ballot aggregation method (explained or not) even if they are among the losers, i.e. the did not vote for a winning project [48]. By considering all 24 combinations of winner vs. loser voters who are satisfied vs. unsatisfied (see Figure 3), we have a rigorous full factorial design to assess the legitimacy of ballot aggregation methods: which method voters prefer (Question MB.4 and MB.7 in Table LABEL:table:entry-voting_input), find fairer (Question MB.6 and MB.9) and more convenient (MB.10).
@@
Proxies of satisfied vs. unsatisfied voters. We use different satisfaction proxies based on questions of the exit survey in Table LABEL:table:entry-digital (OB.1). In particular, we assess whether voters are satisfied with the voting outcome of City Idea (OB.1.1), whether they accept this outcome (OB.1.2) and whether they feel the outcome accurately represents the will of Aarau citizens (OB.1.7). Based on 11-point likert-scale responses, we split the populations two quartiles, the first one for the satisfied voters and the second one for the unsatisfied ones.
@@ 4.3 Explaining shifts between ballot aggregation methods
In this section, we present a causal analysis of how personal traits influence shifts in individual voter preferences for aggregation methods, based on the pre-voting survey, the actual voting process and the post-voting survey.
@@
Prediction algorithms. Prediction models are constructed using tree-based gradient boosting algorithms [49] and neural networks with three hidden layers [50]. Different groups of personal human traits are assessed for the capacity to causally explain shifts between different ballot aggregation methods. The selection of these algorithms aligns well with how they handle categorical data (e.g. CatBoost [51] and neural networks [52] known to perform well even on smaller datasets). Further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the models are illustrated in Section LABEL:sec:shifts-methods.
Approach for causal explanations. Once predictions are made, we derive the relative importance of the personal traits via a two-step explainable framework. Initially, an ablation study [53] is performed to systematically evaluate the relative predictive contribution of each personal human trait. In the second step, we employ model-agnostic Shapley Additive Explanations (SHAP) [54] to assess the individual contribution of each trait to the prediction performance. More details regarding the performance of the explanatory framework can be found in Tables LABEL:tab:accuracy_resultsLABEL:tab:feature-lossLABEL:tab:trait-survey-map and Figures LABEL:fig:ES_G_G_ESLABEL:fig:ES_GLABEL:fig:Convenience_Explain in Section LABEL:sec:shifts-methods.
@@
EP: Conceptualization, Methodology, Validation, Formal analysis, Investigation, Resources, Writing - Original Draft, Visualization, Supervision, Project administration, Funding acquisition. SM: Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Review & Editing, Visualization. TW: Software, Validation, Investigation, Data Curation, Writing - Review & Editing. JCY: Conceptualization, Methodology. FBH: Software, Validation, Data Curation. RHF: Conceptualization, Methodology, Data Collection, Investigation, Resources, Supervision, Writing - Review & Editing, Project administration, Funding acquisition DH: Conceptualization, Investigation, Resources, Writing - Input to Original Draft, Writing - Review & Editing, Supervision, Project administration, Funding acquisition.
@@ Acknowledgments
This project received support by the SNF NRP77 ‘Digital Transformation’ project "Digital Democracy: Innovations in Decision-making Processes", #407740_187249 and the UKRI Future Leaders Fellowship (MR/W009560/1): ‘Digitally Assisted Collective Governance of Smart City Commons–ARTIO’. The authors would like to thank the representatives from the city of Aarau for supporting and running the "Stadtidee" [translated: "City Idea"] campaign. The authors would also like to thank Abhinav Sharma for supporting the team at University of Leeds to collect the voting data as well as Michael Buess from Demoscope for his support with the surveys. The authors are grateful to Lodewijk Gelauff and Ashish Goel for their invaluable support they provided us to extend the Stanford Participatory Budgeting platform.

### v2

@@ 3 Discussion
Fair voting methods to foster a novel impact creation process. The influence of the chosen voting method on the proposed and selected project ideas can transform the process of creating real-world impact. For instance, novel welfare projects gain momentum, as they appeal to several interest groups, while they may also be easier to implement than costly infrastructural projects [Maharjan2024]. We foresee that, on the long term, proportional voting methods could enable an incremental grass-roots implementation of sustainable development projects over several participatory budgeting campaigns, which could also support a fairer transition to more sustainable future.
Fair voting methods are a better-than-optimal paradigm. Over-simplifying, averaging or overriding people’s preferences while optimizing societies for goals that people may not even explicitly agree on is a significant risk that democracies face in the era of artificial intelligence (AI) run by Big Tech. In this context, the notion of ‘optimal’ solutions is actually misleading and displaced from the reality of people, as demonstrated by the inferiority of popular expensive projects or projects focused on city centers that are typically proposed in participatory budgeting processes under standard majoritarian voting.
@@
In contrast, proportional voting methods work well for pluralistic settings, where voters may not agree on one and the same goal, but decisions are based on collective intelligence. This approach excels when goals are complementary, but is also suited to handle situations of competitive or even incompatible goals, or when significant minorities are involved. Small compromises, such as excluding a popular expensive project, open up opportunities for large collective returns with benefits for a great majority of people. For instance, the inclusion of otherwise under-represented health projects or projects in the city periphery can reduce social tensions and benefit a broader population. The demonstrated approach of combining cumulative voting with the method of equal shares delivers a powerful blueprint for upgrading democracies. It allows societies to achieve:
  • (1)
    a larger number of realized projects within the available budget;
  • (2)
    a greater resourcefulness of novel projects and a multi-faceted impact;
  • (3)
    more participation and inclusion;
  • (4)
    greater transparency and trust in the voting process;
  • (5)
    better representation of citizen preferences and higher satisfaction;
  • (6)
    more (spatial) fairness;
  • (7)
    higher legitimacy thanks to a greater acceptance of the collective outcomes;
  • (8)
    strengthened democratic values of voters.
We call these impacts “better than optimal” as they tend to be a win-win for all, while they are distinguished from what existing optimization and data-driven decision-making approaches typically offer.
@@
Fair voting methods in the era of artificial intelligence. We would like to complete this discussion on a remark addressing the use of data-driven AI systems for decision making. While we consider AI a very powerful technology with numerous promising application areas, how to use it in the context of governing democracies must be studied very carefully. AI safeguards for democracy are critical and yet to be created. Existing decision-making methods supported by AI solutions show alarming vulnerabilities and are still lacking fairness and legitimacy [Saetra2022, Helbing2015, Koster2022]. For instance, voting methods with a deficit on fairness can exacerbate biases and inconsistencies by unconsolidated use of large language models [Majumdar2024]. So far, it is not clear how to address some existing AI risks. Therefore, alternative approaches such as the one presented here are urgently needed.
@@ 4 Methods
We illustrate the designed participatory budgeting campaign of “City Idea”, the evaluation approach of legitimacy and the models for causal explanations for the shifts between ballot aggregation methods before and after voting.
4.1 The participatory budgeting process of “City Idea”: Field test design
Table 1 outlines the participatory budgeting campaign of “City Idea”. It consists of six phases over the passage of nine months. An entry and exit survey were conducted before and after voting respectively. The entry survey was launched via physical invitation letters to all citizens sent by city council (10,000 A5 flyers [CityIdeaReport2025]). This resulted in 3592 respondents. Some of these filled in their email to voluntarily participate in the exit survey and this resulted in 808 respondents. The surveys included questions about demographics and political opinion (Table LABEL:table:entry-demographics-LABEL:table:exit-demographics), engagement and digital skills (Table LABEL:table:entry-digital-LABEL:table:exit-digital), the “City Idea” voting (Table LABEL:table:entry-city_idea-LABEL:table:exit-city_idea), the voting outcome preference (Table LABEL:table:entry-voting_outcome-LABEL:table:exit-voting_outcome) and the voting input and aggregation method (Table LABEL:table:entry-voting_input-LABEL:table:exit-voting_input). Survey questions encode variables, which are observed for changes before and after voting. The overall study has received ethical approval from University of Fribourg (#2021-680).
Table 1: The rigorously designed participatory budgeting campaign of “City Idea” with 6 phases over a period of 9 months in 2023. Participants of the entry survey, voting and exit survey are linked using an anonymous identifier (hash) of the citizens’ social security number. This allows the study of several choice variables before and after voting.
@@ 6. Exit survey July September 808 participants
Project ideation, deliberation and campaigning. Citizens proposed more than 161 project ideas out of which 33 projects are selected to put for voting [CityIdeaReport2025]. Projects are eliminated or similar project ideas are combined to strengthen their chances to win. Project ideation involved support and guidance with planning, implementation, economic and legal feasibility by a municipal advisory board of 7 people, a core team of 3 people and ambassadors. The city council distributed posters and flyers to various municipal institutions and locations. It also organized two press releases that resulted in 17 media reports. Moreover, the interactive information campaign ‘know your projects’ was launched in the city that engaged 73 participants in 1682 random pairwise comparisons of projects (see Table LABEL:tab:pairwise_winners for further information) via the Stanford Participatory Budgeting platform  [Wellings2023]. Awareness about the method of equal shares and what projects are expected to win was raised, emphasizing the challenge of electing projects that cover a large share of the budget. Section LABEL:sec:equal-shares-explanation provides more information about how the method of equal shares was communicated and explained to citizens.
Voting. The city of Aarau has a total of 22,593 citizens, out of which 17,428 are Swiss citizens. A total of 1703 voters participated in the "City Idea" campaign, out of which 1330 are citizens of Aarau. This also allows individuals to have a voice, who are not living in Aarau, but have a connection to the city, such as workers and broader family members. Children over 15 years old could also vote. Voting was digital, using the open-source Stanford Participatory Budgeting platform [Wellings2023]. However, voters with low digital literacy were supported in virtual voting centers around the city. The voting platform was extended to support cumulative voting by distributing 10 points over at least 3 projects [SPBStadtidee2023, Wellings2023]. Voters are securely verified in a privacy-preserving way by providing a unique (hash) code that was generated by the city council using the social security number of each citizen (AHV card). The same platform was used for project campaigning by using the pairwise project comparisons feature. The winning projects were calculated using the method of equal shares and the Add1U completion method implemented within the Pabutools (the ones of Świecie and Wieliczka used the Add1 completion method, while Add1U would yield 1 and 3 more winners) [pabutools]. Voting outcomes with the utilitarian completion method were also calculated and assessed in the exit survey (Table LABEL:table:exit-voting_outcome).
@@ 4.2 Evaluation of legitimacy
To assess the legitimacy of ballot aggregation methods, we rely on well-established theory of political legitimacy [Weatherford1992] and social choice [Arrow2010, Brandt2016]. Among the three notions of input (voting participation), throughput (the voting process) and output (voting outcome) legitimacy [Scharpf1999, Schmidt2013], we mainly focus in Section 2.3 on output legitimacy. However, we also assess the choice of ballot aggregation method before and after voting with equal shares as well as the effect of explaining equal shares, both capturing throughput legitimacy.
As such, by legitimacy we refer to voters who remain satisfied and accept the voting outcome of a ballot aggregation method (explained or not) even if they are among the losers, i.e. the did not vote for a winning project [Hanggli2024]. By considering all 24 combinations of winner vs. loser voters who are satisfied vs. unsatisfied (see Figure 3), we have a rigorous full factorial design to assess the legitimacy of ballot aggregation methods: which method voters prefer (Question MB.4 and MB.7 in Table LABEL:table:entry-voting_input), find fairer (Question MB.6 and MB.9) and more convenient (MB.10).
@@
Proxies of satisfied vs. unsatisfied voters. We use different satisfaction proxies based on questions of the exit survey in Table LABEL:table:entry-digital (OB.1). In particular, we assess whether voters are satisfied with the voting outcome of “City Idea” (OB.1.1), whether they accept this outcome (OB.1.2) and whether they feel the outcome accurately represents the will of Aarau citizens (OB.1.7). Based on 11-point likert-scale responses, we split the populations two quartiles, the first one for the satisfied voters and the second one for the unsatisfied ones.
@@ 4.3 Explaining shifts between ballot aggregation methods
In this section, we present a causal analysis of how personal traits influence shifts in individual voter preferences for ballot aggregation methods, based on the pre-voting survey, the actual voting process and the post-voting survey.
@@
Prediction algorithms. Prediction models are constructed using tree-based gradient boosting algorithms [prokhorenkova2018catboost] and neural networks with three hidden layers [singh2019study]. Different groups of personal human traits are assessed for the capacity to causally explain shifts between different ballot aggregation methods. The selection of these algorithms aligns well with how they handle categorical data (e.g. CatBoost [hancock2020catboost] and neural networks [wu2018development] known to perform well even on smaller datasets). Further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the models are illustrated in Section LABEL:sec:shifts-methods.
Approach for causal explanations. Once predictions are made, we derive the relative importance of the personal traits via a two-step explainable framework. Initially, an ablation study [mohseni2021multidisciplinary] is performed to systematically evaluate the relative predictive contribution of each personal human trait. In the second step, we employ model-agnostic Shapley Additive Explanations (SHAP) [nohara2019explanation] to assess the individual contribution of each trait to the prediction performance. More details regarding the performance of the explanatory framework can be found in Tables LABEL:tab:accuracy_resultsLABEL:tab:feature-lossLABEL:tab:trait-survey-map and Figures LABEL:fig:ES_G_G_ESLABEL:fig:ES_GLABEL:fig:Convenience_Explain in Section LABEL:sec:shifts-methods.
@@
EP: Conceptualization, Methodology, Validation, Formal analysis, Investigation, Resources, Writing - Original Draft, Visualization, Supervision, Project administration, Funding acquisition. SM: Software, Validation, Formal analysis, Investigation, Data Curation, Writing - Review & Editing, Visualization. TW: Software, Validation, Investigation, Data Curation, Writing - Review & Editing. JCY: Conceptualization, Methodology. FBH: Software, Validation, Data Curation. RHF: Conceptualization, Methodology, Data Collection, Investigation, Resources, Supervision, Writing - Review & Editing, Project lead and administration, Funding acquisition DH: Conceptualization, Investigation, Resources, Writing - Input to Original Draft, Writing - Review & Editing, Supervision, Project administration, Funding acquisition.
@@ Acknowledgments
This project received support by the SNF NRP77 ‘Digital Transformation’ project “Digital Democracy: Innovations in Decision-making Processes”, #407740_187249 and the UKRI Future Leaders Fellowship (MR/W009560/1): “Digitally Assisted Collective Governance of Smart City Commons–ARTIO”. The authors would like to thank the representatives from the city of Aarau for supporting and running the “Stadtidee” [translated: "City Idea"] campaign. The authors would also like to thank Abhinav Sharma for supporting the team at University of Leeds to collect the voting data, Sajan Maharjan for his support with the Pabulib data, as well as Michael Buess from Demoscope for his support with the surveys. The authors are grateful to Lodewijk Gelauff and Ashish Goel for their invaluable support they provided us to extend the Stanford Participatory Budgeting platform.